COMPLETED: Planned Migration: UKCloud Shared Hosting

Following our acquisition of UKCloud customers, we are in the process of migrating the shared hosting services to our own datacentre and servers.

This process is currently underway and the main bulk of data transfers and updates will be completed by 12:00 BST 25th March 2012.  Any domains using and will be migrated automatically.  Any domains using external nameservers will need to be updated manually – customers should contact support (support at othellotech dot net) or for more information.

COMPLETED: Planned Maintenance – LINX (Extreme LAN) Renumbering

As part of the expansion of LINX (the London Internet eXchange), we are required to change IP addressing and configuration with peering partners.

Saturday 17th March 09:00 onwards.

This work is not expected to cause any issues, although the path traffic is routed along will change during the works.  Othello has more than sufficient connectivity to handle the re-routed traffic during the works.

COMPLETED: Planned Maintenance: Hosting : CPanel Upgrades


RAM Updages on, and


Thursday 23/Feb 20:00 – 21:00

Each server will be offline for less than 10 minutes.


Clients with sites and/or email hosted on, and


Replacing (increasing) server RAM.

COMPLETED: Interservers Server Upgrades and Server Migrations


Interservers Server Upgrades and Server Migrations


Mon 20/Feb through Mon 28/Feb


Clients on Interservers legacy hardware


Migration of UK clients to new UK servers Updates of Server O/S and Control Panel Installation of new Softaculus Script installers.

Clients who have their domain registered with Othello will not experience any downtime, and all changes will be automatic
Clients who have their domain/nameservers elsewhere will be contacted to update their nameservers

COMPLETED: Power Failure at Othello DC 2HQ London

15:25 – There has been a major power failure in the E14 9xx postcode, affecting supply to our DataCentre at Harbour Quay (2HQ).

Automated Transfer Switches have switched over to our secondary powerfeed, and the generators moved to hot status until the incident is resolved by the power company.

No customer or Othello systems have been affected – all critical and client services are fed from our site-wide 2N+1 Uninterruptable Power Supplies (UPS) fed by 2 diverse power systems, backed up by 2N Generators.

COMPLETED: Planned Maintenance/Upgrade – Client Portal

We will be upgrading the Client Portal ( on 24/01/2012 between 18:00 and 19:00 UTC.

During the upgrade existing clients will not be able to access the client portal features such as searching for viewing invoices etc.  Access to individual hosting control panels etc will still be available directly as advised in your welcome email. The upgrade will also prevent new clients from signing up,  checking domain availability etc.


The client portal upgrade has been completed if you experience any problems please raise a support ticket.

COMPLETED: CentralNIC Planned Maintenance

CentralNIC have advised us of their annual database maintenance period.  This is scheduled for 09:00 (GMT/UTC) Sunday January 29 2012.  They did not provide a duration, however we expect the service to be considered at-risk until at least 13:00.  During this period it will not be possible to register, renew or update (contacts or nameservers) CentralNIC based domain names.  Customers who attempt changes will be presented with error messages, renewal and registrations will be queued and manually processed by the Domains Team. (Common CentralNIC TLDs are,, etc.)  The domains will continue to resolve correctly throughout the maintenance window.

Final Update (January 29 2012 – 09:18): This maintenance has been completed.  All CentralNIC functions are available again.

COMPLETED: Planned Maintenance/Upgrade – PHP 5.3 (cPanel Servers)

Commencing 23rd January 2012 we will begin upgrading cPanel  Servers to PHP 5.3.  This will have an impact on scripts that have not been coded to PHP standards or rely on outdated behaviours.

For most pre-packaged applications this should have no effect at all – with the exception of some upgrade choices becoming available.  If your previously working scripts fail after the upgrade, please contact Support via the normal channels (support -at- / to see if we can provide any assistance.  Please note however, that all customer scripts are their responsibility. has full details of current standards and all the details of changes between 5.2 and 5.3.

cPanel1 ( will be upgraded Monday 23rd January 17:00 (approx. GMT).
cPanel2 ( will be upgraded Tuesday 24th January 17:00 (approx. GMT).
cPanel3 ( will be upgraded Wednesday 25th January 17:00 (approx. GMT).
cPanel5 ( will be upgraded Thursday 26th January 17:00 (approx. GMT).
cPanel6 ( will be upgraded Friday 27th January 17:00 (approx. GMT).

These upgrades may be postponed or cancelled depending on the levels of issues encountered with customer scripts.  Please check the status blog ( on the date of your server’s upgrade for final confirmation.

Final Update (27 January 2012 20:59) – All upgrades are now completed.

COMPLETED: Emergency Maintenance : Hosting : Plesk1 (Duck)

Server Plesk1 (Duck) has again suffered a RAID array degredation.  We have therefore treated the hardware as suspect and provisioned a replacement.  User data is currently being migrated to the new hardware, services will be unavailable until this process completes.  We apologise for any inconvenience, but believe this to be the prudent course of action given recent events with this hardware.

ETA for services / Next Update: 12:00 noon 17/December/2011

COMPLETED: Emergency Maintenance : Hosting : Plesk1 (Duck)

Server Plesk1 (Duck) has failed with a corrupt RAID array.

DataCentre technicians are onsite investigating if it can be repaired and returned to service.

If the current server is unrecoverable a new plesk server will be deployed and all sites/emails etc restored from backups.

ETA for server availability / next update 12:30 16/December/2011

Update 1:
We apologise for the lateness of the update; repairs are continuing to progress, data recovery is at approximately 10%.
Next update: 15:00 16/December/2011

Update 2:
This server has now been restored to service.  We apologise for the extended downtime, this machine is being scheduled for replacement in the near future.  E-mail will begin to be delivered over the next few hours, however this depends on the sending server’s retry schedule.
If you have any further issues, please contact support via the usual channels.